OpenStudy (anonymous):

x + 9/11 = 8/11 is the answer -1 show work

OpenStudy (ash2326):

\[x+\frac{9}{11}=\frac{8}{11}\] subtracting both sides by 9/11 \[x=\frac{8}{11}-\frac{9}{11}\] \[x=-\frac{1}{11}\] answer is x=\(-\frac{1}{11}\)

OpenStudy (anonymous):

maybe this is \[\frac{x+9}{11}=\frac{8}{11}\] in which case \[x+9=8\] and \[x=-1\]